OpenStudy (anonymous):

Simplify the expression. 4(6 – 3) A. 7 B. 12 C. 21 D. 24 Please Help (:

OpenStudy (anonymous):

i keep getting 0.75

OpenStudy (anonymous):

You use the distribution method which would make it: 4 x 6 - 4 x 3

OpenStudy (anonymous):

:/ wow thats different but okay :)

OpenStudy (anonymous):

To distribute, you take the number in front and multiply it to the numbers in the parenthesis.

OpenStudy (anonymous):

okay i start from the left frist right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

It would be 24 - 12

OpenStudy (anonymous):

so 12?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Okay thanks soo much

OpenStudy (anonymous):

Glad to be of service.

OpenStudy (anonymous):

:)

OpenStudy (phi):

4(6 – 3) that means 4 times (6-3) order of operations: do stuff in parens first: 6-3 = 3 4*3=12 you can also "distribute the 4" which means 4*6 - 4*3 now order of operations: do multiplies first, then add/subtract so 24 -12 = 12 same answer, a little more work.